The length of a rectangle is 4 times the width. The perimeter is 50 cm. Find the length and width. Can you show me how to do this, maybe by elimination?
Answer by TutorDelphia(193)

l=4w
2l+2w=50 (this is from the perimeter formula 2l+2w=p
You mentioned you wanted to do this using the elimination method. Let's get both equations in good form for that.
l=4w
subtract 4w from both sides
l-4w=0
2l+2w=50
divide both sides by 2
l+w=25
minus
l-4w=0
5w=25 (minus -4w gives us +4w)
w=5
l=20